a number is five less than another number the difference between twice the larger and three times the smaller number is six find the cube of the smaller number
Answer by ankor@dixie-net.com(22740)

let x = "a number
let y = "another number"
:
Write an equation for each statement:
:
"a number is five less than another number"
x = y - 5 (x is the smaller number)
or
y = (x+5)
:
"the difference between twice the larger and three times the smaller number is six"
2y - 3x = 6
replace y with (x+5)
2(x+5) - 3x = 6
2x + 10 - 3x = 6
2x - 3x = 6 - 10
-x = -4
x = 4 is the smaller number
:
Find the cube of the smaller number
4^2 = 64
:
:
Check solution in the statement
"the difference between twice the larger and three times the smaller number is six"
y = 4 + 5 = 9
2(9) - 3(4) =
18 - 12 = 6